# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v2

@@ II.3 Connection to the Born-Oppenheimer framework
If we define the electronic Hamiltonian $\hat{H}^{\mathrm{e}}(\mathbf{Q})$ as the first two terms of Eq. (3) at a fixed nuclear position $\mathbf{Q}$, then we can find an electronic Hamiltonian matrix $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ with elements $H_{\mathbf{mn}}^{\mathrm{e}}(\mathbf{Q})=\langle\mathbf{m}|\hat{H}^{\mathrm{e}}(\mathbf{Q})|\mathbf{n}\rangle_{\mathrm{e}}$, where only electronic degrees of freedom are integrated. The Hamiltonian matrix can be transformed by a unitary matrix $\mathbf{W}(\mathbf{Q})$ to give $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})=\mathbf{W}(\mathbf{Q})\mathbf{H}^{\mathrm{e}}(\mathbf{Q})\mathbf{W}^{\dagger}(\mathbf{Q})$. When $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})$ is a fully diagonal matrix, we arrive at the exact (full configuration interaction, FCI) solution to the electronic structure problem. In practice, full diagonalization is intractable on classical computers, so the eigenvalues of a submatrix of $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ may be found instead. Iterative diagonalization is typically used to further reduce the cost, and the number of eigenvalues found, $N_{\mathrm{BO}}$, can be less than the rank of the (sub)matrix. The BO electronic states are the corresponding eigenfunctions [Fig. 1(d)], and molecular dynamics can be expressed by the Born-Huang expansion as nuclear dynamics on multiple BO states:
@@
Without spin-orbit coupling, the maximum number of BO electronic states is equal to the number of configuration state functions (CSFs, i.e. spin-adapted linear combinations of Slater determinants) with a fixed electron count and spin, $N_{\mathrm{CSF}}$, which we will discuss in more detail in Sec III. The summation over the BO states in Eq. (11) is exact within the given orbital basis set only when $N_{\mathrm{BO}}=N_{\mathrm{CSF}}$. The relation between the Born-Huang expansion and our ansatz [Eq. (2)] becomes clear via the relation $\sum_{j}^{N_{\mathrm{CSF}}}\sum_{\mathbf{v}^{\prime}}T_{j\mathbf{v}^{\prime}}(t)(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=C_{\mathbf{v}\mathbf{n}}(t)$, where $(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=\int\mathrm{d}\mathbf{Q}\,\mathbf{v}^{*}(\mathbf{Q})W_{j\mathbf{n}}(\mathbf{Q})\mathbf{v}^{\prime}(\mathbf{Q})$. However, $N_{\mathrm{CSF}}$ scales rapidly with the orbital basis set size whereas $N_{\mathrm{BO}}$ must be kept small for practical simulations. Therefore, the GBOA is employed ($N_{\mathrm{BO}}<N_{\mathrm{CSF}}$) to reduce the computational cost of the BO state calculations and time-propagation of the molecular vibronic state [Fig. 1(e)]. In contrast, our ansatz [Eq. (2)] is exact within the given vibronic basis by definition.
@@ II.4 Numerical test
We show a proof-of-principle demonstration of our analog simulation approach using the one-dimensional, two-electron Shin-Metiu model [62, 63] and its implementation on a trapped ion quantum simulator, which has been used for the quantum simulation of chemical dynamics for several systems [41, 43, 44, 64]. The model consists of two electrons, two fixed ions, and one moving ion between the fixed ions displaced by a distance $L=5.4$ a.u. [Fig. 3(a)], where the origin is set to the middle point of the two fixed ions. The computational details for all simulations performed in this section, including the parameters for the model Hamiltonian, are given in the Supplemental Material [46].
We construct the electronic Fock space with four spin orbitals $\{|\phi_{i}\rangle\}$ using the spin-up and spin-down configurations of two diabatic spatial orbitals, $\eta_{a}$ and $\eta_{b}$ in Figs. 3(b) and 3(c) respectively, i.e. $\varphi_{1}=\varphi_{3}=\eta_{a}$, and $\varphi_{2}=\varphi_{4}=\eta_{b}$. As a result, a total of four ions are needed in the experimental trapped ion setup for the simulation, with only two of the 12 available motional modes needed for the simulation. The diabatic orbitals were obtained by numerically integrating two adiabatic orbitals from the one-electron Shin-Metiu Hamiltonian [62] subject to the diabatization condition. The resulting orbitals are localized around the left and right fixed ions, but delocalize slightly as the moving ion approaches the fixed ion [Figs. 3(b) and 3(c)]. We replaced the Coulomb potential for $V_{\mathrm{nn}}$ with a harmonic potential to simplify the Hamiltonian and to confine the spatial extent of the wavefunction. The one- and two-electron integrals are truncated at first order in the nuclear position: $v(\hat{Q})\approx v_{0}+v_{1}\hat{Q}$, where $\hat{Q}=(\hat{b}^{\dagger}+\hat{b})/\sqrt{2\omega}$ with ladder operators $\hat{b}^{\dagger}$ and $\hat{b}$ of our single mode with frequency $\omega$ of the harmonic nuclear repulsion potential. With these approximations, the cMQB Hamiltonian can be written as
@@
$N_{\mathrm{CSF}}=3$ is the size of complete electronic basis of singlet electronic states with two electrons in two spatial orbitals. We have excluded the ground BO state from the dynamics with the GBOA ($N_{\mathrm{BO}}=2$), which has relatively small NACs with the first excited state (see Supplemental Material [46] for densities of electronic states in the BO framework). Within a short simulation time (56.1 a.u.), the GBOA density shows a spurious electron transfer from the left fixed ion to the right [Fig. 3(g)] where the subsequent density deviates even further [Fig. 3(j)]. The electron transfer at $t=56.1$ a.u. does not occur in the exact simulation [Fig. 3(e)] due to vibronic coupling to the ground BO state. This demonstrates how the GBOA may result in inaccurate molecular vibronic dynamics when a molecule reaches an unanticipated NAC region coupled to a BO state outside of the truncated BO basis. On the other hand, the Trotterized cMQB time-evolution reproduces the exact density with a converged Trotter step of $\Delta t=5.6$ a.u [Figs. 3(f) and 3(i)] with the fidelity $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}>0.95$, where the convergence of fidelity with respect to the Trotter step is reported in the Supplemental Material [46].
Figure 4: Comparison of noise effects due to motional decoherence with different Trotter step sizes for cMQB simulations. (a) Time-evolution of the fractional occupation number of $\phi_{1}$. (b) Time-evolution of the fidelity of the molecular wavefunctions: $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}$. The blue, red, and green lines represent the noise simulation results with $\Delta t=5.6$, $9.6$, and $16.8$ a.u., respectively, where the dashed lines represent the corresponding closed system simulation results and the black line represents the exact closed system simulation result without Trotterization.
@@
Figure 4(b) confirms the trade-off between Trotter error the and noise effect. The fidelity with $\Delta t=9.6$ a.u. is greater than the fidelity with $\Delta t=16.8$ a.u. due to the smaller Trotter error. However, the fidelity decreases again when the Trotter step is further reduced to $\Delta t=5.6$ a.u. due to the contribution from the digital part in the molecular scaling factor [Eqs. (9) and (10)] for the noise. Therefore, the Trotter step should be chosen carefully considering both the Trotter error and the noise.
@@ III Discussion
We first compare the resource and computational costs of our pre-BO analog cMQB simulation approach with those of existing classical and quantum algorithms, where the advantages of our approach are highlighted in Fig. 5.
The number of qubits required for our approach is equal to the number of spin orbitals, $N_{\mathrm{o}}$. On the other hand, the equivalent number of electronic states in the BO framework corresponds to the exact Born-Huang expansion limit ($N_{\mathrm{BO}}=N_{\mathrm{CSF}}$). Therefore, the classical resources scale proportionally to $N_{\mathrm{CSF}}$ for a fixed spin-multiplicity, which can be calculated according to the Weyl’s formula for the corresponding total spin angular momentum $S$:
$\begin{split}N_{\mathrm{CSF}}(S,N_{\mathrm{o}},N_{\mathrm{e}})&=\\ \frac{2S+1}{N_{\mathrm{o}}/2+1}&\binom{N_{\mathrm{o}}/2+1}{N_{\mathrm{e}}/2-S}\binom{N_{\mathrm{o}}/2+1}{N_{\mathrm{e}}/2+S+1}.\end{split}$ (13)
Here, we discuss singlet states ($S=0$). Most ground-state electron configurations of neutral organic molecules are singlets since they have an even number of electrons and a lack of orbital degeneracy. In general, the classical resources scale as $O(N_{\mathrm{o}}^{-2}(m(m-1)^{1/m-1})^{N_{\mathrm{o}}})$ for $N_{\mathrm{o}}=mN_{\mathrm{e}}$ ($m>1$) at the large $N_{\mathrm{o}}$ limit. The number of singlet CSFs is minimum ($O(N_{\mathrm{o}}^{2})$) when $N_{\mathrm{e}}=2$ or $N_{\mathrm{e}}=N_{\mathrm{o}}-2$, and maximum ($O(2^{N_{\mathrm{o}}}/N_{\mathrm{o}}^{2})$) when $2N_{\mathrm{e}}=N_{\mathrm{o}}$ for a fixed $N_{\mathrm{o}}$. Practical simulations fall in between these two ranges with a number of orbitals roughly proportional to the number of electrons, with the ratio depending on the basis set. For example, conjugated polyene chains with a double-zeta basis set have $N_{\mathrm{o}}/N_{\mathrm{e}}=6$. In Fig. 5(a) we show classical resource scaling for singlet states in comparison with our approach. The linear scaling of our approach has a clear advantage over classical methods, even over the minimum $N_{\mathrm{CSF}}$ case for $N_{\mathrm{o}}>6$.
For the nuclear degrees of freedom, the classical simulation has an exponential resource scaling of $O(\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}})$ where $N_{\nu,\mathrm{bas}}$ is the size of basis for mode $\nu$. Our encoding on cMQB simulators has a linear scaling $O(N_{\mathrm{mode}})$, which is equal to that of an analog mixed qudit-boson (MQB) mapping [41].
@@
Previous quantum algorithms, including the MQB approach, can potentially have a smaller quantum resource scaling for electronic degrees of freedom than our method: $\log_{2}(N_{\mathrm{CSF}})<N_{\mathrm{o}}$ for the complete CSF basis [38, 39, 41] since the CSF space is a subspace of the spin orbital Fock space. However, the simulation cost $N_{\mathrm{op}}$ would scale as $O(N_{\mathrm{CSF}}^{2})$ for electronic degrees of freedom in existing quantum simulation approaches, whereas the cMQB mapping scales as $O(N_{\mathrm{o}}^{4})$ [Fig. 5(b)]. The GBOA is a valid approximation for some photo-initiated reactions provided that the photon energy absorbed by a system is known and the BO states include all energetically accessible electronic states. Nevertheless, even if the number of BO states can be reduced, the BO states are obtained approximately since an exact pre-calculation would require fully diagonalizing the FCI Hamiltonian on a classical computer at a prohibitive cost ($O(N_{\mathrm{CSF}}^{3})$), whereas the cMQB approach requires only the pre-calculation of electron integrals that scales as $O(N_{\mathrm{o}}^{4})$. For the nuclear degrees of freedom, our approach has an advantage over other digital or hybrid quantum algorithms for molecular quantum dynamics [38, 39] since our approach has the same resource scaling as the MQB mapping [41] due to the direct mapping on the bosonic modes.
@@
The noise effect on the cMQB device depends primarily on the Trotter step size and the number of digital gates used to generate multi-qubit entanglement [Eqs. (9) and (10)]. Relative to the direct effect of motional decoherence on modes representing vibrations, the indirect qubit noise is negligible. The effect of noise is highly specific to the target problem and device; however, the scaling factors for the noise effect suggest desirable improvements in both computational and experimental aspects of the cMQB approach for a near-term quantum advantage: reducing $N_{\mathrm{CNOT}}$ via circuit optimization and reducing $t_{\mathrm{CNOT}}$ in cMQB devices.
From a different perspective, noise can be leveraged as an advantage in analog simulators by enabling the simulation of open quantum systems at minimal cost as demonstrated in previous work [67, 41, 61, 68], rather than treating it as an effect that must be suppressed.
@@
As previously mentioned, our approach can describe intersystem crossing by considering spin-orbit coupling electron integrals and ionization by including additional terms in the Hamiltonian [Eq. (3)] without any additional resources. Such simiulations have the potential to make our approach even more advantageous since the classical and GBOA-based quantum simulations require a greater number of basis states for different spin multiplicities. Based on the form of the vibronic Hamiltonian, our Hamiltonian can be considered as an extended version of electron-phonon coupling models used in the solid state physics field such as the Hubbard-Holstein model [70]. Therefore, our approach can be readily extended to dynamics involving intersystem crossing and ionization, and translated onto electron-phonon coupling dynamics in solids.

### v3

@@ II.3 Connection to the Born-Oppenheimer framework
If we define the electronic Hamiltonian $\hat{H}^{\mathrm{e}}(\mathbf{Q})$ as the first two terms of Eq. (3) at a fixed nuclear position $\mathbf{Q}$, then we can find an electronic Hamiltonian matrix $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ with elements $H_{\mathbf{mn}}^{\mathrm{e}}(\mathbf{Q})=\langle\mathbf{m}|\hat{H}^{\mathrm{e}}(\mathbf{Q})|\mathbf{n}\rangle_{\mathrm{e}}$, where only electronic degrees of freedom are integrated. The Hamiltonian matrix can be transformed by a unitary matrix $\mathbf{W}(\mathbf{Q})$ to give $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})=\mathbf{W}(\mathbf{Q})\mathbf{H}^{\mathrm{e}}(\mathbf{Q})\mathbf{W}^{\dagger}(\mathbf{Q})$. When $\mathbf{E}^{\mathrm{e}}(\mathbf{Q})$ is a fully diagonal matrix, we arrive at the exact (full configuration interaction, FCI) solution to the electronic structure problem. In practice, full diagonalization is intractable on classical computers, so the eigenvalues of a submatrix of $\mathbf{H}^{\mathrm{e}}(\mathbf{Q})$ may be found instead. For example, full diagonalization within an active space is known as complete active space configuration interaction (CASCI). Iterative diagonalization is typically used to further reduce the cost, and the number of eigenvalues found, $N_{\mathrm{BO}}$, can be less than the rank of the (sub)matrix. The BO electronic states are the corresponding eigenfunctions [Fig. 1(d)], and molecular dynamics can be expressed by the Born-Huang expansion as nuclear dynamics on multiple BO states:
@@
Without spin-orbit coupling, the maximum number of BO electronic states is equal to the number of configuration state functions (CSFs, i.e. spin-adapted linear combinations of Slater determinants) with a fixed electron count and spin, $N_{\mathrm{CSF}}$, which we will discuss in more detail in Sec III. The summation over the BO states in Eq. (11) is exact within the given orbital basis set only when $N_{\mathrm{BO}}=N_{\mathrm{CSF}}$. The relation between the Born-Huang expansion and our ansatz [Eq. (2)] becomes clear via the relation $\sum_{j}^{N_{\mathrm{CSF}}}\sum_{\mathbf{v}^{\prime}}T_{j\mathbf{v}^{\prime}}(t)(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=C_{\mathbf{v}\mathbf{n}}(t)$, where $(W_{j\mathbf{n}})_{\mathbf{v}\mathbf{v}^{\prime}}=\int\mathrm{d}\mathbf{Q}\,\mathbf{v}^{*}(\mathbf{Q})W_{j\mathbf{n}}(\mathbf{Q})\mathbf{v}^{\prime}(\mathbf{Q})$. However, $N_{\mathrm{CSF}}$ scales rapidly with the orbital basis set size whereas $N_{\mathrm{BO}}$ must be kept small for practical simulations. Therefore, the GBOA is employed ($N_{\mathrm{BO}}<N_{\mathrm{CSF}}$) to reduce the computational cost of the BO state calculations and time-propagation of the molecular vibronic state [Fig. 1(e)] based on prior knowledge on the system and its dynamics.
@@ II.4 Numerical test
We show a proof-of-principle demonstration of our digital-analog simulation approach using the one-dimensional, two-electron Shin-Metiu model [45, 46] and its implementation on a trapped ion quantum simulator, which has been used for the quantum simulation of chemical dynamics for several systems [41, 43, 44, 66]. The model consists of two electrons, two fixed ions, and one moving ion between the fixed ions displaced by a distance $L=5.4$ a.u. [Fig. 3(a)], where the origin is set to the middle point of the two fixed ions. The model is an extension of the original model system inspired by nonadiabatic charge transfer in processes with polarization of the medium induced by electron movement [45, 67], such as proton-coupled electron transfer. The computational details for all simulations performed in this section, including the parameters for the model Hamiltonian, are given in the Supplemental Material [48].
We construct the electronic Fock space with four spin orbitals $\{|\phi_{i}\rangle\}$ using the spin-up and spin-down configurations of two diabatic spatial orbitals, $\eta_{a}$ and $\eta_{b}$ in Figs. 3(b) and 3(c) respectively, i.e. $\varphi_{1}=\varphi_{3}=\eta_{a}$, and $\varphi_{2}=\varphi_{4}=\eta_{b}$. As a result, a total of four ions are needed in the experimental trapped ion setup for the simulation, with only two of the 12 available motional modes needed for the simulation. The diabatic orbitals were obtained by numerically integrating two adiabatic orbitals from the one-electron Shin-Metiu Hamiltonian [45] subject to the diabatization condition. The resulting orbitals are localized around the left and right fixed ions, but delocalize slightly as the moving ion approaches the fixed ion [Fig. 3(b) and 3(c)]. We replaced the Coulomb potential for $V_{\mathrm{nn}}$ with a harmonic potential to simplify the Hamiltonian and to confine the spatial extent of the wavefunction. The one- and two-electron integrals are truncated at first order in the nuclear position: $v(\hat{Q})\approx v_{0}+v_{1}\hat{Q}$, where $\hat{Q}=(\hat{b}^{\dagger}+\hat{b})/\sqrt{2\omega}$ with ladder operators $\hat{b}^{\dagger}$ and $\hat{b}$ of our single mode with frequency $\omega$ of the harmonic nuclear repulsion potential. With these approximations, the cMQB Hamiltonian can be written as
@@
$N_{\mathrm{CSF}}=3$ is the size of complete electronic basis of singlet electronic states with two electrons in two spatial orbitals. We have excluded the ground BO state from the dynamics with the GBOA ($N_{\mathrm{BO}}=2$), which has relatively small NACs with the first excited state (see Supplemental Material [48] for densities of electronic states in the BO framework). Within a short simulation time (56.1 a.u.), the GBOA density shows a spurious electron transfer from the left fixed ion to the right [Fig. 3(g)] where the subsequent density deviates even further [Fig. 3(j)]. The electron transfer at $t=56.1$ a.u. does not occur in the exact simulation [Fig. 3(e)] due to vibronic coupling to the ground BO state. The inaccuracy of the GBOA in this example is much greater than one would expect for realistic molecules, due to the minimal nature of the model. On the other hand, the Trotterized cMQB time-evolution reproduces the exact density with a converged Trotter step of $\Delta t=5.6$ a.u [Figs. 3(f) and 3(i)] with the fidelity $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}>0.95$, where the convergence of fidelity with respect to the Trotter step is reported in the Supplemental Material [48]. A more detailed discussion on the comparison of our approach with the BO framework is given in Discussion.
Figure 4: Comparison of errors originating from noise effects due to motional decoherence and Trotterization, with different Trotter step sizes for cMQB simulations. (a) Time-evolution of the fractional occupation number of $\phi_{1}$. (b) Time-evolution of the fidelity of the molecular wavefunctions: $|\langle\Psi|\Psi_{\mathrm{exact}}\rangle|^{2}$. The blue, red, and green lines represent the noise simulation results with $\Delta t=5.6$, $9.6$, and $16.8$ a.u., respectively, where the dashed lines represent the corresponding closed system simulation results and the black line represents the exact closed system simulation result without Trotterization.
@@
Fig. 4(b) confirms the trade-off between the Trotter error and the noise effect. The fidelity with $\Delta t=9.6$ a.u. is greater than the fidelity with $\Delta t=16.8$ a.u. due to the smaller Trotter error. However, the fidelity decreases again when the Trotter step is further reduced to $\Delta t=5.6$ a.u. due to the contribution from the digital part in the molecular scaling factor [Eqs. (9) and (10)) for the noise. Therefore, the Trotter step should be chosen carefully considering both the Trotter error and the noise.
@@ III Discussion
We first compare the resource and computational costs of our pre-BO cMQB simulation approach with those of existing classical and quantum algorithms, where the advantages of our approach are highlighted in Fig. 5. Then, we discuss considerations for practical applications including noise effects and Trotter errors, which depend on the device parameters of cMQB simulators.
The number of qubits required for our approach is equal to the number of spin orbitals, $N_{\mathrm{o}}$. On the other hand, the equivalent number of electronic states in the BO framework corresponds to the exact Born-Huang expansion limit ($N_{\mathrm{BO}}=N_{\mathrm{CSF}}$). Even for cases where $N_{\mathrm{BO}}$ can be kept small, the resources required for the classical calculation of the BO states equivalent to our approach (i.e. FCI, or CASCI for a subset of orbitals) scales with $N_{\mathrm{CSF}}$. Therefore, the classical resources of the BO framework equivalent to our approach scale proportionally to $N_{\mathrm{CSF}}$ for a fixed spin-multiplicity, which can be calculated according to the Weyl’s formula for the corresponding total spin angular momentum $S$:
[…]
Here, we discuss singlet states ($S=0$). Most ground-state electron configurations of neutral organic molecules are singlets since they have an even number of electrons and a lack of orbital degeneracy. In general, the classical resources scale as $O(N_{\mathrm{o}}^{-2}(m(m-1)^{1/m-1})^{N_{\mathrm{o}}})$ for $N_{\mathrm{o}}=mN_{\mathrm{e}}$ ($m>1$) at the large $N_{\mathrm{o}}$ limit. The number of singlet CSFs is minimum ($O(N_{\mathrm{o}}^{2})$) when $N_{\mathrm{e}}=2$ or $N_{\mathrm{e}}=N_{\mathrm{o}}-2$, and maximum ($O(2^{N_{\mathrm{o}}}/N_{\mathrm{o}}^{2})$) when $2N_{\mathrm{e}}=N_{\mathrm{o}}$ for a fixed $N_{\mathrm{o}}$. Practical simulations fall in between these two ranges with a number of orbitals roughly proportional to the number of electrons, with the ratio depending on the basis set. For example, conjugated polyene chains with a double-zeta basis set have $N_{\mathrm{o}}/N_{\mathrm{e}}=6$. In Fig. 5(a) we show classical resource scaling for singlet states in comparison with our approach. The linear scaling of our approach has a clear advantage over classical methods, even over the minimum $N_{\mathrm{CSF}}$ case for $N_{\mathrm{o}}>6$. Other electronic structure methods may use a different wavefunction ansatz with lower resource costs, but they are not equivalent to our approach, making their comparison difficult due to the many factors involved. The classical simulation has an exponential resource scaling for nuclear degrees of freedom of $O(\prod_{\nu=1}^{N_{\mathrm{mode}}}N_{\nu,\mathrm{bas}})$ where $N_{\nu,\mathrm{bas}}$ is the size of basis for mode $\nu$, whereas our encoding has a linear scaling $O(N_{\mathrm{mode}})$.
@@
Previous quantum algorithms, including the MQB approach, can potentially have a smaller quantum resource scaling for electronic degrees of freedom than our method: $\log_{2}(N_{\mathrm{CSF}})<N_{\mathrm{o}}$ for the complete CSF basis [38, 39, 41] since the CSF space is a subspace of the spin orbital Fock space. However, the simulation cost $N_{\mathrm{op}}$ would scale as $O(N_{\mathrm{CSF}}^{2})$ for electronic degrees of freedom in existing quantum simulation approaches, whereas the cMQB mapping scales as $O(N_{\mathrm{o}}^{4})$ [Fig. 5(b)]. For the nuclear degrees of freedom, our approach has an advantage over other digital or hybrid quantum algorithms for molecular quantum dynamics [38, 39] since our approach has the same resource scaling as the MQB mapping [41] due to the direct mapping on the bosonic modes.
The GBOA reduces $N_{\mathrm{op}}$ to $O(N_{\mathrm{BO}}^{2})$, and its use is justified where a limited number of BO electronic states are accessible at different nuclear geometries. Nevertheless, even if the number of BO states can be reduced, the BO states may be obtained approximately since an exact pre-calculation would require fully diagonalizing the FCI or CASCI Hamiltonian on a classical computer at a prohibitive cost ($O(N_{\mathrm{CSF}}^{3})$), whereas the cMQB approach requires only the pre-calculation of electron integrals that scales as $O(N_{\mathrm{o}}^{4})$. Highly accurate electronic structure methods have a cost scaling greater than our approach ($O(N_{\mathrm{o}}^{6\sim 8})$[70], whereas methods with a lower accuracy require careful characterization. Furthermore, the scalar coupling terms $\sum_{\nu}\langle\Phi_{j}^{\mathrm{BO}}|\tfrac{\partial^{2}}{\partial Q^{2}_{\nu}}|\Phi_{k}^{\mathrm{BO}}\rangle$ are impractical to calculate for most electronic structure methods and are often neglected, whereas the analogous terms are easily included in our pre-BO approach.
@@
The noise effect on the cMQB device depends primarily on the Trotter step size and the number of digital gates used to generate multi-qubit entanglement (Eqs. (9) and (10)), which has a trade-off with the Trotter error. Relative to the direct effect of motional decoherence on modes representing vibrations, the indirect qubit noise is negligible. The effect of noise is highly specific to the target problem and device; however, the scaling factors for the noise effect suggest desirable improvements in both computational and experimental aspects of the cMQB approach for a near-term quantum advantage: reducing $N_{\mathrm{CNOT}}$ via circuit optimization and reducing $t_{\mathrm{CNOT}}$ in cMQB devices.
From a different perspective, noise can be leveraged as an advantage in analog simulators by enabling the simulation of open quantum systems at minimal cost as demonstrated in previous work [71, 41, 65, 72], rather than treating it as an effect that must be suppressed. For example, our model problem can be interpreted as a nonadiabatic charge transfer process occurring in solution under the influence of solvent interactions, or within a crystal at an interface with a substrate.
Considering the quantum resource requirements and noise effects, our approach can be implemented on current and near-term hardware. Early applications are well suited to molecules with a small active space, particularly rigid molecules whose nuclear dependence is well-described by low-order Taylor expansions. These molecules are also the target of efficient classical algorithms, but our approach has advantages for the accuracy by including all terms in the vibronic Hamiltonian for the choice of orbital basis. However, practical advantage should be examined carefully by assessing the achievable efficiency and accuracy in the BO framework and the effect of errors in near-term cMQB devices.
For long term applications, we note that the range of orbital energies in a molecule are typically much greater than the range of energetically accessible BO electronic states. The Trotter step size $\Delta t_{\mathrm{mol}}$ decreases proportional to the orbital energy range. This in turn decreases the scaling factor $F$ due to the minimum experimental time step, leading to longer simulation times and thus proportionally greater noise. To reduce the energy range, we can choose a set of active orbitals at the cost of reducing simulation accuracy. Because our approach is more strongly limited by the orbital energies rather than the number of orbitals, we expect it to accommodate larger active spaces than classical computing approaches. For this reason, our approach will be particularly advantageous for strongly correlated systems, which can have large numbers of nearly degenerate orbitals, leading to extremely large numbers of CSFs and intractable BO-based calculations.
Nuclear quantum effects, such as proton tunneling, are naturally included in our approach. Such effects are essential for describing many chemical processes, especially in biological systems [73]. Our choice of model system represents two static moieties (fixed ions) coupled to two electrons and a moving ion (e.g. a proton). The time evolution resulted in a superposition of the moving ion between left and right positions, demonstrating the importance of the quantum description. Furthermore, the results for the model system demonstrate how small changes in the electronic description can result in large differences in the dynamics. These results demonstrate that our approach will enable accurate and efficient simulations of the quantum effects in photochemical dynamics and proton-coupled electron transfer.
@@
As previously mentioned, our approach can describe intersystem crossing by considering spin-orbit coupling electron integrals and ionization by including additional terms in the Hamiltonian [Eq. (3)] and incorporating the corresponding operations in the cMQB simulator without any additional quantum resources. Such simulations have the potential to make our approach even more advantageous since the classical and quantum simulations in the BO framework require a greater number of basis states and the corresponding electronic structure calculations for different spin multiplicities. Based on the form of the vibronic Hamiltonian, our Hamiltonian can be considered as an extended version of electron-phonon coupling models used in the solid state physics field such as the Hubbard-Holstein model [75]. Therefore, our approach can be readily extended to dynamics involving intersystem crossing and ionization, and translated onto electron-phonon coupling dynamics in solids.